OpenStudy (anonymous):

The expected value of two random variables, E[XY] can be defined as an inner space defined by those variables as . The lecture notes I am looking at say that the norm associated with this inner space is ||Y||=^1/2. Shouldn't it be that the associated norm is ||XY||=^1/2? I thought the norm of an inner space was defined as the square root of that inner space (as in the pythagorean theorem gives the "norm" of the triangle.) Thanks.

OpenStudy (anonymous):

XY is not an element of your inner product space, but Y is

OpenStudy (anonymous):

Let take an easy inner product space, for example the plane X=(a,b), Y=(c,d), then <X,Y>= a c + bd ||Y||^2 =c^2 + d^2 =<Y,Y>

OpenStudy (anonymous):

So that means that there is another norm associated with the same space? Specifically ||X||=<X,X?>1/2 right?

OpenStudy (anonymous):

I guess that must be true since X and Y are symmetric. I guess I was thrown off partially by the notes saying it was "THE associated norm" in stead of "An associated norm.

OpenStudy (anonymous):

For every scalar product space there is a norm associated with it